OpenStudy (anonymous):

Easton simplified the following expression: (x^2y^6z^5)(x^4y^5z^3) If he writes his answer in the form of x^a y^b z^c, what is the value of b, the exponent on y?

OpenStudy (anonymous):

\[x^6 y^11 z^8\]

OpenStudy (anonymous):

(x^2y^6z^5)(x^4y^5z^3) = x^6y^11z^8 so b = 11

OpenStudy (anonymous):

(x^2 y^6 z^5) (x^4 y^5 z^3) -> y^6 * y^5 = y^11 Thus b = 11